# Approximate predictive control barrier function for discrete-time systems

### v1

@@ 6.2 Nonlinear system
where the length of the car is $l=0.09$m and the sampling time is $T_{s}=0.02$s. The states are given by the $x$- and $y$-position of the car, the heading angle $\psi$, the velocity $v$ and the steering angle $\delta$, i.e. $x=\left[p_{x},p_{y},\psi,v,\delta\right]^{\top}\in\mathbb{R}^{5}$ and the inputs are given by an acceleration and the steering change, resulting in $u\in\mathbb{R}^{2}$. The input constraints are given by $-5\leq u_{0}\leq 2$ and $-2.8\leq u_{1}\leq 2.8$ and the steering lies within $-0.75\leq\delta\leq 0.75$. The car operates in a cluttered environment with 9 obstacles, which are defined by $(p_{x}-o_{x,i})^{2}+(p_{y}-o_{y,i})^{2}\geq r_{i}^{2}$ for each obstacle $i$ with varying centre position $(o_{x,i},o_{y,i})$ and radius $r_{i}$. Finally, there are no constraints on the other states. For compactness of $\mathcal{D}$ and $\mathcal{S}$, they can be assumed to be large enough that they remain inactive during operation.
@@
[…]
@@
This implementation aims to reach the opposite side of the initial condition of the area, while keeping the velocity constant at $v_{\textup{ref}}=3\textup{ms}^{-1}$. Notably, the optimisation problem is not aware of the obstacles, such that obstacle avoidance is purely a result of the safety filter. The closed-loop results for $5$ initial conditions over $70$ time steps are shown in Figure 3, where the system successfully avoids the obstacles in the cluttered environment, even at the high initial speed of $3\textup{ms}^{-1}$. The solve time for the neural network inference for safety verification, i.e. lines 4 and 5 in Algorithm 1, take on average $0.26$ms and solving the two optimisation problems takes on average $37$ms. Comparatively, the PCBF algorithm in [12] requires on average $500$ms to solve the two predictive optimisation problems. The computational speedup is therefore a factor of $13$ in the case of overwriting the inputs and $1900$ when the proposed input is deemed safe. Finally, we provide the magnitude of the safety interventions for one of the trajectories in Figure 4. It can be seen that the safety filter not only overrides the steer change before an obstacle is hit, but also the acceleration such that the car is able to navigate around the obstacles safely.
Figure 3: Closed-loop trajectories of the simulated car dynamics, where multiple initial conditions with a velocity of $3$m/s and an initial steering angle of $\delta=0$ are shown (car). The performance controller is given by a naive MPC implementation which is not aware of the obstacles, such that the distance to the opposite side of the area is minimised. States where the safety filter intervened are marked in red.
Figure 4: Magnitude of safety filter interventions of the steer change (top) and the acceleration (bottom) for the trajectory with initial position $(-1.8,-1.0)$.
@@ 7 CONCLUSION
We propose a computationally efficient safety filter formulation which relies on approximating the optimal value function of a predictive control method. This function encodes the information regarding safety in terms of state constraint satisfaction with respect to the entire prediction horizon of the predictive optimisation problem and enhances the safe set and domain of attraction of a terminal control barrier function. By using the explicit approximation, a horizon independent optimisation problem can be formulated with a single decrease constraint of the control barrier function. We advance the theoretical properties with respect to exogeneous disturbances of the original algorithm in [12] and translate the results to input-to-state stability of the proposed algorithm with respect to approximation errors. Finally, we propose a negative continuous extension of the predictive control barrier function in order to mitigate approximation errors within the safe set. The theoretical properties are validated in an illustrative linear example and the practicality of a computationally advantageous safety filter is shown in a miniature race car simulation.
In order to achieve a computationally efficient sampling of the extended PCBF $h$ in (21), we propose a modification of the geometric sampling method presented in [41] in Algorithm 2. The geometric sampling method allows using previously computed solutions to the predictive optimisation problems for the computation of $h$ as a warmstart as well as storing information if the previous sample on a line search was within $\mathcal{S}_{\mathrm{PB}}$ or $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, reducing the number of times two optimisation problems have to be solved.
Algorithm 2 Sampling $h$
@@
12:           Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (22) with warmstart $\mathbf{o}_{i-1}$
@@
15:              Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (10) with warmstart $\mathbf{o}_{i-1}$
@@
20:           $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h(x_{i}))$
@@
22:           Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (10) with warmstart $\mathbf{o}_{i-1}$
@@
25:              Obtain $\mathbf{o}_{i}$, $h(x_{i})$ from (22) with warmstart $\mathbf{o}_{i-1}$
@@
30:           $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h(x_{i}))$

### v2

@@ 6.2 Nonlinear System
where the length of the car is l=0.09l=0.09m and the sampling time is $T_{s}=0.02$s. The states are given by the $x$- and $y$-position of the car, the heading angle $\psi$, the velocity $v$ and the steering angle $\delta$, i.e. $x=\left[p_{x},p_{y},\psi,v,\delta\right]^{\top}\in\mathbb{R}^{5}$ and the inputs are given by an acceleration and the steering change, resulting in $u\in\mathbb{R}^{2}$. The input constraints are given by $5\leq u_{0}\leq 2$ and $2.8\leq u_{1}\leq 2.8$ and the steering lies within $0.75\leq\delta\leq 0.75$. The car operates in a cluttered environment with 9 obstacles, which are defined by $(p_{x}-o_{x,i})^{2}+(p_{y}-o_{y,i})^{2}\geq r_{i}^{2}$ for each obstacle $i$ with varying centre position $(o_{x,i},o_{y,i})$ and radius $r_{i}$. Finally, there are no constraints on the other states. For compactness of $\mathcal{D}$ and $\mathcal{S}$, they can be assumed to be large enough that they remain inactive during operation.
@@
The training and validation data sets $\mathbb{D}_{\textup{tr}}=\{(x_{i},h(x_{i})\}$ and $\mathbb{D}_{\textup{val}}$ are collected through the geometric sampler as described in Algorithm 2, with different initial seed and goal points such that the data is independent. The data is collected using $48$ cores on the ETH Euler cluster. The number of collected training and validation points is approximately $1.2\textup{e}7$ and $6\textup{e}6$, with a computation time of 2020h and 1616h, respectively. Given the fact that the approximation error for the negative values within $\mathcal{S}_{\mathrm{PB}}$ and for large values of $h_{\mathrm{PB}}$ have a lower impact on the stability guarantees as also discussed in Remark 5.23, we preprocess the data by applying a cubic root to the function output, i.e. $y_{i}=\sqrt[3]{h(x_{i})}$. Thereby, negative values in $\mathcal{S}_{\mathrm{PB}}$ remain negative and large positive values are decreased, while values close to $\partial\mathcal{S}_{\mathrm{PB}}$ are increased, such that approximation errors in this region incur a higher loss. We train a neural network with $5$ hidden layers of $32$ softplus activation functions on the preprocessed data using $20$ epochs, a learning rate of $2\textup{e}-2$ and a batch size of $128$ using a Huber loss function. The training time was 2.22.2h and the final validation error is given by approximately $1\textup{e}-2$. Note that in principle the loss could be modified to prevent false positives in terms of safety certification of the inputs, as similarly done in [48], by encouraging the regressor to upper bound the true PCBF, but the approximation in this example was sufficiently accurate.
@@
This implementation aims to the reflection of the initial position through the origin, while keeping the velocity constant at $v_{\textup{ref}}=3\textup{ms}^{-1}$. Notably, the optimisation problem is not aware of the obstacles, such that obstacle avoidance is purely a result of the safety filter. The closed-loop results for $5$ initial conditions over $70$ time steps are shown in Figure 4, where the system successfully avoids the obstacles in the cluttered environment, even at the high initial speed of $3\textup{ms}^{-1}$. The solve time for the neural network inference for safety verification, i.e. lines 4 and 5 in Algorithm 1, take on average 0.260.26ms and solving the two optimisation problems takes on average 3737ms. Comparatively, the PCBF algorithm in [12] requires on average 500500ms to solve the two predictive optimisation problems. The computational speedup is therefore a factor of $13$ in the case of overwriting the inputs and $1900$ when the proposed input is deemed safe. Finally, we provide the magnitude of the safety interventions for one of the trajectories in Figure 5. It can be seen that the safety filter not only overrides the steer change before an obstacle is hit, but also the acceleration such that the car is able to navigate around the obstacles safely.
Figure 4: Closed-loop trajectories of the simulated car dynamics, where multiple initial conditions with a velocity of 33m/s and an initial steering angle of $\delta=0$ are shown (car). The performance controller is given by a naive MPC implementation which is not aware of the obstacles, such that the distance to the opposite side of the area is minimised. States where the safety filter intervened are marked in red.
Figure 5: Magnitude of safety filter interventions of the steer change (top) and the acceleration (bottom) for the trajectory with initial position (1.8,1.0)(-1.8,-1.0).
@@ 7 CONCLUSION
We extend initial results in [13], introducing an approximation of the PCBF in a CBF-based safety filter. We advance the theoretical analysis with respect to exogenous disturbances of the algorithms in [12, 13] and propose a negative continuous extension of the predictive control barrier function to mitigate approximation errors within the safe set. The theoretical properties are validated in an illustrative linear example and the practicality of a computationally advantageous safety filter is shown in a miniature race car simulation. Future work could involve integrating the PCBF framework in a reinforcement learning setting similar to [50] to scale the training to possibly higher state dimensions as well as including contextual information of the environment.
In order to achieve a computationally efficient sampling of the extended PCBF $h_{\textup{CE}}$ in (13), we propose a modification of the geometric sampling method presented in [41] in Algorithm 2. The geometric sampling method allows using previously computed solutions to the predictive optimisation problems for the computation of $h_{\textup{CE}}$ as a warmstart as well as storing information if the previous sample on a line search was within $\mathcal{S}_{\mathrm{PB}}$ or $\mathcal{D}_{\mathrm{PB}}\setminus\mathcal{S}_{\mathrm{PB}}$, reducing the number of times two optimisation problems have to be solved.
Algorithm 2 Sampling $h_{\textup{CE}}$
@@
12:    Obtain $\mathbf{o}_{i}$, $h_{\textup{CE}}(x_{i})$ from (14) with warmstart $\mathbf{o}_{i-1}$
@@
15:     Obtain $\mathbf{o}_{i}$, $h_{\textup{CE}}(x_{i})$ from (9) with warmstart $\mathbf{o}_{i-1}$
@@
20:    $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h_{\textup{CE}}(x_{i}))$
@@
22:    Obtain $\mathbf{o}_{i}$, $h_{\textup{CE}}(x_{i})$ from (9) with warmstart $\mathbf{o}_{i-1}$
@@
25:     Obtain $\mathbf{o}_{i}$, $h_{\textup{CE}}(x_{i})$ from (14) with warmstart $\mathbf{o}_{i-1}$
@@
30:    $\mathbb{D}\leftarrow\mathbb{D}\cup(x_{i},h_{\textup{CE}}(x_{i}))$